For the inverse demand curve
P=60-0.5q, find the elasticity of demand at P=10
1
Expert's answer
2019-12-03T10:03:45-0500

"P = 60 - 0.5Q" , "Q = 120 - 2P","P = 10" ,

Q = 120 - 2*10 = 100.

elasticity of demand, Ed = a*P/Q,

Ed = -2*10/100 = -0.2,

Hence, demand is inelastic.